OpenStudy (anonymous):

consider the following: f(x) = tan((pi/x)/2) find the x-values at which f is not continuous. (From the calculatorI know that at -3,-1,1,3, it shows up as error, but when I put them as the answer, it shows up as incorrect)

OpenStudy (jdoe0001):

if allowed, try graphing it, check the vertical asymptotes

OpenStudy (anonymous):

well the first way to approach this is think when is tan undefined because there is where it will not be continuous. we know that as pi/2 and -pi/2 from [0,pi) so set it up as such (pi/x)/2= pi/2 and (pi/x)/2=-pi/2

OpenStudy (anonymous):

-3 and 3 wouldn't give you the answer either pi/3/2 = pi/6